OpenStudy (anonymous):

How do i use the distance formula to show that the equation of the unit circle is given by x^2 + y^2 =1?

OpenStudy (anonymous):

you want the distance from a point on the circle to the origin to be 1. So you the distance formula with the points (x,y) and (0,0), and set it equal to 1.

OpenStudy (anonymous):

Thats correct. Now play around with that for a bit. It will lead to the equation you want.

OpenStudy (anonymous):

\[\sqrt{(x-0)^2+(y-0)^2=1}\]

OpenStudy (anonymous):

and then \[\sqrt{x^2+y^2}=1\] or no?